OpenStudy (anonymous):

A person tries to heat up her bath water by adding 5.0 L of water at 80 degrees celcious to 60 L of water at 30 degrees celcious. What is the final temperature of the water? I got 41.33 C

OpenStudy (anonymous):

The heat lost by the hot water = heat gained by the cool water 5(80-T) = 60(T-30) 400 - 5 T = 60 T - 1800 2200 = 65 T T = 33.8